• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Gomez

    Log entry time 09:16:12 on October 26,1998

    Entry number 3697

    keyword=Hadron dipole ps - comm failure

    During the change to new P0 = 2.796 from 3.028, hadron dipole power supply
    failed to respond. Problem was traced to serial communication card which behaved
    like the communication buffers were full. A soft reset (by login into the IOC and
    performing a reboot or ctrl-X) was performed but it did not clear the problem.
    A hard reset was then performed (green button - middle room). Problem cleared.
    Problem must have been due to an internal register of the serial communication
    card being stuck.